# Beyond Monolingual Assumptions: A Survey of Code-Switched NLP in the Era of Large Language Models

### v1

@@ 1 Introduction
Code-switching (CSW), the alternation between two or more languages within a single utterance or discourse is a pervasive feature of multilingual communication worldwide (Poplack, 1988). With the rise of digital platforms, code-switched text has become ubiquitous across social media and online communication (Molina et al., 2016; Singh and Solorio, 2017), challenging NLP systems built on monolingual assumptions. In India, 26% of the population is bilingual and 7% trilingual (Chandramouli, 2011), with over 250 million speakers engaging in code-switched discourse, yet ASR systems trained on monolingual speech exhibit 30–50% higher word error rates on CSW data (Singh et al., 2025). Multilingual NLU models also suffer up to a 15% drop in semantic accuracy (Winata et al., 2021), leaving bilingual users underserved (Cihan et al., 2022). Figure 1 depicts intra- and inter-sentential code-mixing across multiple language pairs, emphasizing the linguistic variability that NLP systems must navigate.
@@
  • We provide the first comprehensive survey of CSW research in the LLM era, covering NLU and NLG tasks, datasets & benchmarks across diverse language pairs, real-world applications, and key architectural innovations in CSW-aware language processing.
@@
  • We identify key gaps in LLM-based CSW research, including limited low-resource support, script diversity challenges, weak cross-lingual transfer, and the absence of unified evaluation frameworks. Future progress depends on inclusive datasets, equitable models, and fair metrics.
@@ 2 Pre-LLM-Era works
Code-Mixed Language AnalyticsTasksDatasets & ResourcesModel Training & AdaptationEvaluation & BenchmarkingMulti- & Cross-Modal ApplicationsNatural LanguageUnderstanding TasksNatural LanguageGeneration TasksDatasetsFrameworks& ToolkitsFine-tuning ApproachesPre-trainingApproachesPost-trainingApproachesBenchmarksEvaluation MetricsVision-LanguageProcessingSpeech ProcessingCross-ModalIntegration Sentiment & Emotion Analysis: Hinglish SentiMix (Baroi et al., 2020),Spanglish sentiment analysis (Patwa et al., 2020), EmoMix-3L (Bangla-English-Hindi) (Raihan et al., 2023b), multi-label emotion (Ameer et al., 2022; Dowlagar and Mamidi, 2021b) Syntactic Analysis: SyMCoM metrics (Kodali et al., 2022b), dependency parsing (Ghosh et al., 2019; Özateş et al., 2022) Named Entity Recognition: Cross-lingual NER (Zhou et al., 2022; Malmasi et al., 2022b; S and Shrivastava, 2022), entity boundary detection (Kannada-English pair)(Nandigam et al., 2022),(Hindi-English pair)(Singh et al., 2018a) POS Tagging: XLM-RoBERTa fine-tuning & transformer POS tagging(Absar, 2025),joint learning(Kartik et al., 2024; Chatterjee et al., 2022; Dowlagar and Mamidi, 2021c) Language Identification: word-level LID (Lambebo Tonja et al., 2022), Boundary detection (Sterner and Teufel, 2023), MaskLID (Kargaran et al., 2024), ECT (Kuwanto et al., 2024), LCT (Iliescu et al., 2021), Hope speech detection Hossain et al. (2021); Dave et al. (2021), Offensive language detection (Chakravarthi et al., 2021; Goswami et al., 2023; Jayanthi and Gupta, 2021; Balouchzahi et al., 2021a) Natural Language Inference: Cross-lingual NLI tasks (Chakravarthy et al., 2020; Shankar et al., 2024; Khanuja et al., 2020a; Prasad et al., 2021) Question Answering: MLQA(Lewis et al., 2020; Kumar et al., 2022b), XQuAD evaluation frameworks (Wu et al., 2025b), unified CSQA benchmark (Gupta et al., 2018; Chandu et al., 2018a) Intent Classification: Multilingual intent prediction and slot filling (Krishnan et al., 2021; Srivastava and Singh, 2020; Raihan et al., 2023a) Cross-lingual Transfer: Zero-shot transfer learning (Wu et al., 2023; Do et al., 2024; Lai et al., 2021; Whitehouse et al., 2022) Machine Translation: CS-to-English MT (Dhar et al., 2022; Kang, 2024; Borisov et al., 2025), pivot translation strategies (Huzaifah et al., 2024) Code-Mixed Text Generation: Synthetic data creation (Potter and Yuan, 2024; Salaam et al., 2022; Liu et al., 2022), LLM prompting (Kuwanto et al., 2024; Pratapa et al., 2018a) Text Summarization: Abstractive and extractive approaches (Mehnaz et al., 2021; Arora et al., 2023; Yoo et al., 2025) Transliteration: Indic language transliteration (Anand and Kumar, 2022; Vavre et al., 2022; Dowlagar and Mamidi, 2021b), script normalization (Nag et al., 2024) Dialogue Generation: Conversational agents(Hindi-English pair) (Agarwal et al., 2023; Liu et al., 2022; Kumar and Chakraborty, 2024) Multilingual Coverage: SwitchLingua (Xie et al., 2025),Megaverse (Ahuja et al., 2024),GLUECoS (Khanuja et al., 2020b) Low-Resource Coverage: Kazakh-Russian (KRCS) (Borisov et al., 2025; Balouchzahi et al., 2022; Goloburda et al., 2025),African (Ogunremi et al., 2023),Kanglish (Balouchzahi et al., 2022),Basque-Spanish (Heredia et al., 2025a),Moroccan Darija-French (Samih and Maier, 2016) Synthetic Data Generation: LLM-generated (Kuwanto et al., 2024), POS tagging Kartik et al. (2024), quality estimation (Hamed et al., 2025; Sravani and Mamidi, 2023; Srivastava and Singh, 2022a) Annotation Frameworks: COMMENTATOR (Sheth et al., 2024), ToxVidLM (Maity et al., 2024),Corpus (Kochar et al., 2024) Synthetic Data Generation Toolkits: Based on linguistic theories (Kuwanto et al., 2024; Gautam et al., 2021b), LLM-based generation (Potter and Yuan, 2024) Task-Adaptive Pre-Training (TAPT): Domain-specific adaptation and in-domain unlabeled data (Gururangan et al., 2020; Nayak and Joshi, 2022; Goswami et al., 2023) Specialized Code-Mixed Models: HingBERT (Nayak and Joshi, 2022), CoMix (Arora et al., 2023),COMMIT (Lee et al., 2024),SetFit Pannerselvam et al. (2024) Cross-lingual Alignment: COMMIT approach (Lee et al., 2024), continual learning (Zhang et al., 2021), embedding alignment (Lample et al., 2018) Instruction Tuning: (Lee et al., 2024; Wang et al., 2024; Zhang et al., 2025b) Multi-task fine-tuning: Joint training across CS tasks (Aguilar et al., 2020; Zhang et al., 2025b) Task-specific fine-tuning: Progressive complexity (Agro et al., 2025), switching point focus, easier-to-harder training (Yoo et al., 2025) Parameter Efficient Methods: QLoRA (Srivastava, 2025), knowledge distillation (Sidibomma et al., 2025) Reinforcement Learning for CSW Adaptation: RL with reward models & policy gradient methods (Palivela et al., 2025; Zhang et al., 2025c) Zero-, One- and Few-shot Learning: robustness improvement (Winata et al., 2021; Zhang et al., 2023),in-context learning Shankar et al. (2024); Held et al. (2023) Instance-Based Prompting: PRO-CS technique (Bansal et al., 2022; Udawatta et al., 2024; Deroy and Maity, 2025) Comprehensive Benchmarks: GLUECoS (Khanuja et al., 2020b), LinCE (Aguilar et al., 2020), MEGAVERSE (22 languages) (Ahuja et al., 2024), DravidianCodeMix (Chakravarthi et al., 2022) Domain-Specific Corpora: Social media (Chakravarthi et al., 2022), agricultural (Olaleye et al., 2025), conversational (Budzianowski et al., 2018) Task-Specific Metrics: chrF for MT (Popović, 2015), WER for ASR (Kadaoui et al., 2024), human evaluation protocols (Shimorina and Belz, 2022) CS-Specific Metrics: CMI (Chatterjere et al., 2020), SyMCoM (Kodali et al., 2022b) Traditional Metrics: F1-score (Raihan et al., 2023b), BLEU (Jasim et al., 2020), accuracy measures (Sterner and Teufel, 2025b) Quality Assessment: PhoBLEU (Arora et al., 2023), semantic preservation scores (Reimers and Gurevych, 2020), naturalness evaluation (Liu et al., 2021) Intrinsic Evaluation: Language modeling perplexity (Chen and Goodman, 1996), grammaticality scoring (Warstadt et al., 2019) Audio-Visual Recognition: CI-AVSR Cantonese (Dai et al., 2022), multimodal speech processing Raj Khan et al. (2021), multimodal speech processing Speech Translation: Direct speech-to-text translation (P S V N et al., 2025), End speech to speech (Weller et al., 2022), ASR and MoE system (Zhang et al., 2025a) End-to-End ASR: Code-switched speech recognition (Kronis et al., 2024), monolingual model fine-tuning (Bansal et al., 2022) Document Processing: Multilingual OCR (Dereza et al., 2024), code-switched document analysis (Do et al., 2024) Multimodal VQA: Knowledge distillation approaches (Raj Khan et al., 2021), code-mixed visual question answering (Ahuja et al., 2024) Phonetic Processing: Transliteration modules (Weisberg Mitelman et al., 2024), phonetic embedding for ASR (Chi and Bell, 2022) Multimodal Fusion: Audio-visual-text systems for hate speech (Srivastava, 2025), cross-modal alignment for emotion (Wu et al., 2025a)
@@ 3.1 Natural Language Understanding Tasks
LLM-based approaches improve CSW POS tagging via pre-trained models, multi-task learning, and transfer strategies. mBERT embeddings capture Arabic–English CSW nuances, and character-CNNs with mBERT further boost Hinglish POS (Sabty et al., 2020; Aguilar and Solorio, 2020). Joint LID–POS transformer models and switch-point–biased self-training improve accuracy on Telugu–English data (Dowlagar and Mamidi, 2021c; Chopra et al., 2021). Bilingual pretraining on GLUECoS enhances POS (Winata et al., 2021; Prasad et al., 2021). PACMAN supports fine-tuning, while PRO-CS and CoMix leverages prompts for improved POS tagging and generation of code-switched text (Chatterjee et al., 2022; Kumar et al., 2022b; Bansal et al., 2022; Arora et al., 2023). XLM-R fine-tuned with the S-index achieves state-of-the-art on Hinglish POS improves prformance (Absar, 2025).
@@ 3.2 Natural Language Generation Tasks
A detailed discussion of remaining NLU and NLG tasks is provided in Appendix §A.1 and §A.2.
@@ 9 Conclusion
CSW research has undergone a major transformation with the rise of LLMs. From early rule-based and statistical methods to multilingual pretraining, fine-tuning, and instruction-based adaptation, the field has steadily moved toward more unified and scalable approaches. This survey shows that while significant progress has been made in LID, tagging, translation, and generation for high-resource pairs, major gaps persist for low-resource and culturally diverse contexts. The future of CSW research lies in building inclusive datasets, improving evaluation protocols, and designing socially grounded systems that reflect the realities of multilingual communication.
@@ Appendix A Methodology
We began by defining a set of search keywords targeting three core dimensions: code-mixing/code-switching, multilingual NLP, and large language models. To ensure broad linguistic coverage, the search included major bilingual and multilingual language pairs documented in prior works and repositories (e.g., CoVoSwitch, GLUECoS, LinCE). Using these keywords, we queried the ACL Anthology, arXiv and Semantic Scholar databases via their APIs, with a search cutoff date of October 2025, consistent with ACL Rolling Review’s recency guidelines. This process initially retrieved  500 papers.
@@
Duplicate entries were removed using DOIs and titles, prioritizing peer-reviewed sources. The remaining papers were manually screened for relevance. A study was included if it addressed code-mixing or code-switching within any NLP task, or explored multilingual and LLM-based adaptation methods. This screening resulted in a refined set of 304 papers, covering both pre-LLM and LLM-era research.
@@ A.5 Fine-tuning Approaches
The image 3 shows a ChatGPT hallucination where a query for vacation countries results in a list of Indian states and union territories. The response also exhibits incorrect grammar, demonstrating the model’s limitations in handling nuanced, code-mixed queries.
The image 4 shows a failure case from Perplexity AI. When prompted in a code-mixed query for Goan temples specifically destroyed by the Portuguese, the model disregards this critical constraint and instead provides information on existing ancient temples.
@@
The image 5 shows an illustration of hallucination in the open-source model GLM-4.6. When prompted to perform morphophonemic (sandhi) analysis of the Bengali word Ohoroho (“always”) in Bengali-Romanized Hindi script, the model misinterprets the query and generates an incorrect response.

### v3

@@ 1 Introduction
Code-switching (CSW), the alternation between two or more languages within a single utterance or discourse, is a pervasive feature of multilingual communication worldwide (Poplack, 1988). With the rise of digital platforms, code-switched text has become ubiquitous across social media and online communication (Molina et al., 2016; Singh and Solorio, 2017), challenging NLP systems built on monolingual assumptions. In India, 26% of the population is bilingual and 7% trilingual (Chandramouli, 2011), with over 250 million speakers engaging in code-switched discourse, yet ASR systems trained on monolingual speech exhibit 30–50% higher word error rates on CSW data (Singh et al., 2025). Multilingual NLU models also suffer up to a 15% drop in semantic accuracy (Winata et al., 2021), leaving bilingual users under-served (Cihan et al., 2022). Figure 1 depicts intra- and inter-sentential code-mixing across multiple language pairs, emphasizing the linguistic variability that NLP systems must navigate.
@@
  • We provide the first comprehensive survey of CSW research in the LLM era, covering NLU and NLG tasks, datasets and benchmarks across diverse language pairs, real-world applications, and key architectural innovations.
@@
  • We identify key gaps in LLM-based CSW research, including limited low-resource languages support, script diversity challenges, weak cross-lingual transfer, and the absence of unified evaluation frameworks. Future progress depends on inclusive datasets, equitable models, and fair metrics.
@@ 2 Pre-LLM-Era works
Code-Mixed Language AnalyticsTasksDatasets & ResourcesModel Training & AdaptationEvaluation & BenchmarkingMulti- & Cross-Modal ApplicationsNatural LanguageUnderstanding TasksNatural LanguageGeneration TasksDatasetsFrameworks& ToolkitsFine-tuning ApproachesPre-trainingApproachesPost-trainingApproachesBenchmarksEvaluation MetricsVision-LanguageProcessingSpeech ProcessingCross-ModalIntegration Sentiment & Emotion Analysis: Hinglish SentiMix (Baroi et al., 2020),Spanglish sentiment analysis (Patwa et al., 2020), EmoMix-3L (Bangla-English-Hindi) (Raihan et al., 2023b), multi-label emotion (Ameer et al., 2022; Dowlagar and Mamidi, 2021b) Syntactic Analysis: SyMCoM metrics (Kodali et al., 2022b), dependency parsing (Ghosh et al., 2019; Özateş et al., 2022) Named Entity Recognition: Cross-lingual NER (Zhou et al., 2022; Malmasi et al., 2022b; S and Shrivastava, 2022), entity boundary detection (Kannada-English pair)(Nandigam et al., 2022),(Hindi-English pair)(Singh et al., 2018a) POS Tagging: XLM-RoBERTa fine-tuning & transformer POS tagging(Absar, 2025),joint learning(Kartik et al., 2024; Chatterjee et al., 2022; Dowlagar and Mamidi, 2021c) Language Identification: word-level LID (Lambebo Tonja et al., 2022), Boundary detection (Sterner and Teufel, 2023), MaskLID (Kargaran et al., 2024), ECT (Kuwanto et al., 2024), LCT (Iliescu et al., 2021), Hope speech detection Hossain et al. (2021); Dave et al. (2021), Offensive language detection (Chakravarthi et al., 2021; Goswami et al., 2023; Jayanthi and Gupta, 2021; Balouchzahi et al., 2021a) Natural Language Inference: Cross-lingual NLI tasks (Chakravarthy et al., 2020; Shankar et al., 2024; Khanuja et al., 2020a; Prasad et al., 2021) Question Answering: MLQA(Lewis et al., 2020; Kumar et al., 2022b), XQuAD evaluation frameworks (Wu et al., 2025b), unified CSQA benchmark (Gupta et al., 2018; Chandu et al., 2018a) Intent Classification: Multilingual intent prediction and slot filling (Krishnan et al., 2021; Srivastava and Singh, 2020; Raihan et al., 2023a) Cross-lingual Transfer: Zero-shot transfer learning (Wu et al., 2023; Do et al., 2024; Lai et al., 2021; Whitehouse et al., 2022) Machine Translation: CS-to-English MT (Dhar et al., 2022; Kang, 2024; Borisov et al., 2025), pivot translation strategies (Huzaifah et al., 2024) Code-Mixed Text Generation: Synthetic data creation (Potter and Yuan, 2024; Salaam et al., 2022; Liu et al., 2022), LLM prompting (Kuwanto et al., 2024; Pratapa et al., 2018a) Text Summarization: Abstractive and extractive approaches (Mehnaz et al., 2021; Arora et al., 2023; Yoo et al., 2025) Transliteration: Indic language transliteration (Anand and Kumar, 2022; Vavre et al., 2022; Dowlagar and Mamidi, 2021b), script normalization (Nag et al., 2024) Dialogue Generation: Conversational agents (Hinglish pair) (Agarwal et al., 2023; Liu et al., 2022; Kumar and Chakraborty, 2024) Multilingual Coverage: SwitchLingua (Xie et al., 2025), Megaverse (Ahuja et al., 2024), GLUECoS (Khanuja et al., 2020b) Low-Resource Coverage: Kazakh-Russian (KRCS) (Borisov et al., 2025; Balouchzahi et al., 2022; Goloburda et al., 2025), African (Ogunremi et al., 2023), Kanglish (Balouchzahi et al., 2022), Basque-Spanish (Heredia et al., 2025a), Moroccan Darija-French (Samih and Maier, 2016) Synthetic Data Generation: LLM-generated (Kuwanto et al., 2024), POS tagging Kartik et al. (2024), quality estimation (Hamed et al., 2025; Sravani and Mamidi, 2023; Srivastava and Singh, 2022a) Annotation Frameworks: COMMENTATOR (Sheth et al., 2024), ToxVidLM (Maity et al., 2024), Corpus (Kochar et al., 2024) Synthetic Data Generation Toolkits: Based on linguistic theories (Kuwanto et al., 2024; Gautam et al., 2021b), LLM-based generation (Potter and Yuan, 2024) Task-Adaptive Pre-Training (TAPT): Domain-specific adaptation and in-domain unlabeled data (Gururangan et al., 2020; Nayak and Joshi, 2022; Goswami et al., 2023) Specialized Code-Mixed Models: HingBERT (Nayak and Joshi, 2022), CoMix (Arora et al., 2023), COMMIT (Lee et al., 2024), SetFit Pannerselvam et al. (2024) Cross-lingual Alignment: COMMIT approach (Lee et al., 2024), continual learning (Zhang et al., 2021), embedding alignment (Lample et al., 2018) Instruction Tuning: (Lee et al., 2024; Wang et al., 2024; Zhang et al., 2025b) Multi-task fine-tuning: Joint training across CS tasks (Aguilar et al., 2020; Zhang et al., 2025b) Task-specific fine-tuning: Progressive complexity (Agro et al., 2025), switching point focus, easier-to-harder training (Yoo et al., 2025) Parameter Efficient Methods: QLoRA (Srivastava, 2025), knowledge distillation (Sidibomma et al., 2025) Reinforcement Learning for CSW Adaptation: RL with reward models & policy gradient methods (Palivela et al., 2025; Zhang et al., 2025c) Zero-, One- and Few-shot Learning: robustness improvement (Winata et al., 2021; Zhang et al., 2023),in-context learning Shankar et al. (2024); Held et al. (2023) Instance-Based Prompting: PRO-CS technique (Bansal et al., 2022; Udawatta et al., 2024; Deroy and Maity, 2025) Comprehensive Benchmarks: GLUECoS (Khanuja et al., 2020b), LinCE (Aguilar et al., 2020), MEGAVERSE (22 languages) (Ahuja et al., 2024), DravidianCodeMix (Chakravarthi et al., 2022) Domain-Specific Corpora: Social media (Chakravarthi et al., 2022), agricultural (Olaleye et al., 2025), conversational (Budzianowski et al., 2018) Task-Specific Metrics: chrF for MT (Popović, 2015), WER for ASR (Kadaoui et al., 2024), human evaluation protocols (Shimorina and Belz, 2022) CS-Specific Metrics: CMI (Chatterjere et al., 2020), SyMCoM (Kodali et al., 2022b) Traditional Metrics: F1-score (Raihan et al., 2023b), BLEU (Jasim et al., 2020), accuracy measures (Sterner and Teufel, 2025b) Quality Assessment: PhoBLEU (Arora et al., 2023), semantic preservation scores (Reimers and Gurevych, 2020), naturalness evaluation (Liu et al., 2021) Intrinsic Evaluation: Language modeling perplexity (Chen and Goodman, 1996), grammaticality scoring (Warstadt et al., 2019) Audio-Visual Recognition: CI-AVSR Cantonese (Dai et al., 2022), multimodal speech processing Raj Khan et al. (2021), multimodal speech processing Speech Translation: Direct speech-to-text translation (P S V N et al., 2025), End speech to speech (Weller et al., 2022), ASR and MoE system (Zhang et al., 2025a) End-to-End ASR: Code-switched speech recognition (Kronis et al., 2024), monolingual model fine-tuning (Bansal et al., 2022) Document Processing: Multilingual OCR (Dereza et al., 2024), code-switched document analysis (Do et al., 2024) Multimodal VQA: Knowledge distillation approaches (Raj Khan et al., 2021), code-mixed visual question answering (Ahuja et al., 2024) Phonetic Processing: Transliteration modules (Weisberg Mitelman et al., 2024), phonetic embedding for ASR (Chi and Bell, 2022) Multimodal Fusion: Audio-visual-text systems for hate speech (Srivastava, 2025), cross-modal alignment for emotion (Wu et al., 2025a)
@@ 3.1 Natural Language Understanding Tasks
LLM-based approaches improve CSW POS tagging accuracy via pre-trained models, multi-task learning, and transfer strategies. mBERT embeddings capture Arabic–English CSW nuances, and character-CNNs with mBERT further boost Hinglish POS (Sabty et al., 2020; Aguilar and Solorio, 2020). Joint LID–POS transformer models and switch-point–biased self-training improve accuracy on Telugu–English data (Dowlagar and Mamidi, 2021c; Chopra et al., 2021). Bilingual pretraining on GLUECoS enhances POS (Winata et al., 2021; Prasad et al., 2021). PACMAN supports fine-tuning, while PRO-CS and CoMix leverages prompts for improved POS tagging and generation of code-switched text (Chatterjee et al., 2022; Kumar et al., 2022b; Bansal et al., 2022; Arora et al., 2023). XLM-R fine-tuned with the S-index achieves state-of-the-art on Hinglish POS improves performance (Absar, 2025).
@@ 3.2 Natural Language Generation Tasks
A detailed discussion of remaining NLU and NLG tasks is provided in Appendix §A.1 and §A.2, while the corresponding datasets and approaches for NLP tasks are summarized in §A.5 Table 2.
@@ 9 Conclusion
CSW research has undergone a major transformation with the rise of LLMs. From early rule-based and statistical methods to multilingual pretraining, fine-tuning, and instruction-based adaptation, the field has steadily moved toward more unified and scalable approaches. This survey shows that although significant progress has been made in LID, tagging, translation, and generation for high-resource pairs, major gaps persist for low-resource and culturally diverse contexts. The future of CSW research lies in building inclusive datasets, improving evaluation protocols, and designing socially grounded systems that reflect the realities of multilingual communication.
@@ Appendix A Methodology
We began by defining a set of search keywords targeting three core dimensions: code-mixing/code-switching, multilingual NLP, and large language models. To ensure broad linguistic coverage, the search encompassed major bilingual and multilingual language pairs documented in prior research and repositories (e.g., CoVoSwitch, GLUECoS, LinCE). Using these keywords, we queried the ACL Anthology, arXiv and Semantic Scholar databases via their APIs, with a search cutoff date of October 2025, consistent with ACL Rolling Review’s recency guidelines. This process initially retrieved  500 papers.
@@
Duplicate entries were removed using DOIs and titles, prioritizing peer-reviewed sources. The remaining papers were manually screened for relevance. A study was included if it addressed code-mixing or code-switching within any NLP task, or explored multilingual and LLM-based adaptation methods. This screening resulted in a refined set of 308 papers, covering both pre-LLM and LLM-era research.
@@ A.5 Fine-tuning Approaches
Figure 3 shows a ChatGPT hallucination where a query for vacation countries results in a list of Indian states and union territories. The response also exhibits incorrect grammar, demonstrating the model’s limitations in handling nuanced, code-mixed queries.
Figure 4 shows an illustration of hallucination in the open-source model GLM-4.6. When prompted to perform morphophonemic (sandhi) analysis of the Bengali word Ohoroho (“always”) in Bengali-Romanized Hindi script, the model misinterprets the query and generates an incorrect response.
Figure 5 shows a failure case from Perplexity AI. When prompted in a code-mixed query for Goan temples specifically destroyed by the Portuguese, the model disregards this constraint and instead provides information on only existing ancient temples and not the destroyed ones.